### v3

# DeToNATION: Decoupled Torch Network-Aware Training on Interlinked Online Nodes

Abstract
Training large neural network models requires extensive computational resources, often distributed across several nodes and accelerators. Recent findings suggest that it may be sufficient to only exchange the fast moving components of the gradients, while accumulating momentum locally (Decoupled Momentum, or DeMo). However, DeMo assumes that models fit on a single accelerator. We relax this assumption and introduce FlexDeMo, whereby nodes fully shard model parameters locally between different accelerators, while inter-node communication is reduced by synchronizing only fast-moving components instead of the full gradients – resulting in a hybrid sharded data parallel training strategy. We further introduce a framework, denoted as DeToNATION, that generalizes DeMo, FlexDeMo, and other popular distributed training schemes such as DiLoCo – introducing new variations of replication schemes and challenging choices made in DeMo. Our results across language and vision domains show that FlexDeMo attains similar validation loss as hybrid sharded data parallel training employing AdamW and full gradient synchronization, while being substantially faster. FlexDeMo is thus a promising distributed training scheme for the largest machine learning models.
Codehttps://github.com/schneiderkamplab/DeToNATION/
Introduction
Training large deep neural networks (DNNs) induces large amounts of network traffic in the form of gradients that are transmitted between accelerators, typically requiring expensive localized high-throughput network setups on high-performance computing clusters. The network throughput increasingly becomes a bottleneck, as the number of accelerator nodes participating in the training increases and the general network congestion increases. Recent work demonstrates that synchronizing the full optimizer state is not always necessary for state-of-the-art results through decoupling momentum updates and allowing controlled divergence within each rank by carefully controlled inter-accelerator communication termed Decoupled Momentum optimization (DeMo) (Peng, Quesnelle, and Kingma 2024).
DeMo presents a viable strategy for distributed training with reduced gradient communication and, thus, enables distributed data parallel (DDP) training of DNNs with relatively low network bandwidth requirements. However, this strategy comes with several caveats. First, relying on DDP implies the constraint that the DNN model and optimizer states must fit within the memory capacity of each accelerator. This severely limits the applicability for training large models such as state-of-the-art large language models, which commonly do not fit within the memory of a single accelerator. Second, DeMo inherently relies on a distributed gathering operation whose bandwidth requirements scale linearly with the number of accelerators. Third, DeMo leaves quite a lot of open questions regarding the replication-scheme and hyperparameters, such as chunk-size, TopK and employing the sign-function.
In this work, we introduce the Flexible Decoupled Momentum optimizer (FlexDeMo) for combining fully sharded data parallel (FSDP) training with decoupled momentum updates. This optimizer employs a hybrid-sharding strategy, where the model and optimizer states are typically sharded intra-node and replicated between nodes.
Instead of synchronizing the full gradients between the nodes as in extant hybrid sharding strategies, we compress, gather, and decompress selected relevant fast-moving momentum components following the approach that the DeMo optimizer has introduced for DDP training (Peng, Quesnelle, and Kingma 2024). This relaxes constraints regarding the accelerator memory requirements while simultaneously reducing inter-node bandwidth requirements. We further investigate hyper-parameters of DeMo and challenge the fast-moving momenta-based replication scheme by introducing Striding, Random and the prior proposed DiLoCo schemes. This addresses the three caveats of DeMo mentioned above. First, FlexDeMo allows for decoupled momentum training of models that do not fit into the memory of a single accelerator but fit into the combined memory of the accelerators of one node. Second, FlexDeMo replicates between nodes rather than accelerators, effectively reducing the bandwidth requirements of the distributed gathering operation, which now scales linearly with the number of nodes rather than the number of accelerators. Third, we justify our choice of hyper-parameters by conducting experiments.
We validate our results on the T5, ViT and OLMo models in machine translation, image classification and causal language modeling, respectively. We demonstrate that FlexDeMo is faster than, and comparable in terms of validation loss, to both DeMo and extant hybrid sharding strategies that combine intra-node FSDP with full inter-node gradient synchronization. In essence, FlexDeMo enables training of larger DNNs more efficiently across multiple nodes of a cluster or even across geographically-dispersed clusters. We demonstrate that the Random replication-scheme, given the same bandwidth usage significantly outperforms DeMo and other replications schemes in translation, whereas we demonstrate DeMo to be superior in image-classification and causal language modeling. Last, we show that sharing the signed gradients is clearly beneficial.
In summary, our contributions are as follows:
  • The first implementation of a hybrid sharded training strategy combining intra-node FSDP with decoupled momentum optimization (FlexDeMo).
  • Increased efficiency in bandwidth-limited settings, both compared to FSDP with full gradient synchronization and previous work (DeMo, DiLoCo).
  • Comparable model performance compared to extant hybrid sharded training strategies using the AdamW optimizer.
  • Introducing Random, Striding and DiLoCo replication schemes into a new DeToNATION framework.
  • Introducing a decoupled variant of AdamW.
  • An analysis of the critical hyper-parameters TopK, chunk-size and sign providing guidance on its effects on efficiency and model performance.
Related Work
Strategies to both scale and accelerate the training of deep neural networks have been active research areas (Shoeybi et al. 2019; Rajbhandari et al. 2020) for quite some years. The importance of these methods effectiveness is increasing as we train more and more large language models both in and across growing HPC capacities. We will first clarify the terminology and then provide a brief overview of the most important advances in distributed training.
Distributed data parallel (DDP) replicates the model and optimizer states across multiple individual processes, which each handle a subset of the training data. The gradients are averaged from all processes to keep the model weights synchronized. In model parallelism, the model is split across accelerators, consequently increasing the communication overhead as each device computes only the forward and backward passes for its assigned model-parts, requiring the communication of the intermediate activation between devices. Tensor parallelism is similar to model parallelism but even individual tensors are split across devices.
The zero redundancy optimizer (ZeRO) (Rajbhandari et al. 2020), tackling data and model parallelism, has introduced a variety of strategies to train large models, including partitioning optimizer states, gradients, and parameters – all of which are being pulled dynamically to a single accelerator on demand. ZeRO has been integrated in the DeepSpeed library (Rasley et al. 2020). The DeepSpeed team has since iterated on ZeRO with releases of Zero-2 (DeepSpeed Team, Majumder, and Wang 2020) and Zero-3 (DeepSpeed Team 2021). While ZeRO aims to achieve memory-efficiency to enable training of very large models, it also comes with a severe communication overhead. Both DeMo (Peng, Quesnelle, and Kingma 2024) and our proposed FlexDeMo aim to alleviate this communication overhead.
FSDP (Zhao et al. 2023) takes inspiration from ZeRO, but makes a few modifications to the exchange of gradients: ZeRO uses a reduce-scatter operation to distribute the gradients and an all-gather operation for the updated parameters. FSDP instead uses an all-gather operation to recover unsharded parameters and a reduce-scatter operation for the gradients. FSDP also introduces a hybrid sharding strategy, enabling the user to define the set of sharding accelerators and corresponding replicates to allow for a trade-off between memory and throughput via sharding groups and gradient accumulation with or without communication.
Another common strategy is to optimize locally for multiple steps before synchronization (Stich 2019). DiLoCo (Douillard et al. 2023) employs federated averaging to enable training in poorly connected nodes through parallel local optimization and periodic global averaging.
Notably, there are several other advances in accelerating large scale neural network training, such as compressing the gradient by only considering its sign (Bernstein et al. 2018; Jiang et al. 2024) or through low-rank projection of the gradients (Zhao et al. 2024). Although tackling similar challenges, we consider these directions orthogonal to our work, as we are interested in improving the communication overhead regarding optimizer states.
Methods
We first present our hybrid-sharded decoupled momentum optimizer, FlexDeMo, which allows divergent optimizer states across accelerator nodes. Sharding to the accelerators of a node and replicating between the nodes is a reasonable strategy. We introduce the Random and Striding replication schemes, compare with the priorly proposed DiLoCo, and challenge DeMo’s effectiveness both regarding convergence and computation. FlexDeMo can be used to shard and replicate between any sets of accelerators, though, allowing both replicating model shards to multiple subsets of accelerators on one node and distributing different shards across multiple nodes. Thereby, we are relaxing the model memory-constraint by sharding the model and optimizer states across multiple accelerators, typically within one node, it becomes possible to train significantly larger models. DeMo replication is guided by the fast moving components of the momentum, computed via the discrete cosine transform (DCT-II), $n$-strides indices for Striding and, naturally, randomly selected for Random.
One of the main practical limitations of DeMo is that it does not support FSDP or other sharding strategies at all and relies on an all_gather operation across multiple nodes, which is inherently slow, as it requires every training process to exchange and receive gradients from every other instance (see Appendix A). This raises the new research question regarding whether DeMo would be suitable in an sharded setting such as FSDP, as well as investigating if and how its hyperparameters need to be adjusted – in particular, how many fast-moving components to extract and further, if such components is most optimal to synchronize. Here, we set out to overcome these technical hurdles and investigate and validate the efficiency characteristics of such distributed training mechanism.
FlexDeMo
During training, the optimizer states are communicated between the accelerators within the node. Then fast moving components are extracted following the method from DeMo (Peng, Quesnelle, and Kingma 2024). These components are compressed and exchanged between groups of accelerators. This method allows for training larger models, keeping the expensive communication within the groups of accelerators, and minimizing the expensive communication between groups.
The flexible decoupled momentum training strategy we propose can be powered by multiple replication-schemes; The DeMo based Discrete Cosine Transform (DCT) of the momenta, Random, Striding and DiLoCo. The replication of the optimizer states is done between the groups of accelerators instead of between all accelerators and by only communicating the selected components (gradients). This is shown in the Appendix A, where accelerator 0 of node 0 is replicating momentum components to accelerator 0 of node 1, and so on. With this method, there is no replication of data between accelerator 0 in one node and accelerator 1 in another, effectively drastically limiting the amount of data shared between nodes.
While we only sharing the selected gradients (e.g.: fast moving momentum components) between the nodes, and even only between accelerators with corresponding shards, we operate on the full gradients within each group, where the transfer speeds are usually significantly higher. This takes advantage of the typically high bandwidth within nodes, while acknowledging the slower bandwidth between nodes.
To achieve this, we reimplement the stochastic gradient descent (SGD) with momentum optimizer from (Peng, Quesnelle, and Kingma 2024) (denoted DeMo-SGD), introducing a series of changes to support decoupled optimization in FSDP employing intra-node hybrid-sharding. We assume that the model is wrapped as an FSDP object. Both the forward and backward passes must be wrapped in a no_sync context manager, disabling automatic synchronization of gradients $\theta_{t}$ affecting the default all_reduce-functionality, consequently decoupling the momentum $m$ across accelerator-nodes. PyTorch’s Autograd produces the gradient for the whole unsharded grad-parameter accessible in p.grad.
Shown in Appendix A is a schematic representation of the communication. We employ the reduce-scatter operation, averaging and then sharding the computed gradients back to their respective ranks (hence sharded parameter-size) in the sharding-parallel-group. This allow us to work only on the shards in the subsequent operations. We denote the two communication-groups as $S$ and $R$ referring to the sharding-group and replication-group, respectively. Gradients are scattered intra-node locally within $S$ and the fast components, $q$, are communicated inter-node in $R$. We describe the operation of FlexDeMo in Algorithm 1.
The figure in Appendix A shows the communication pattern for standard DeMo within a single node, as well as the communication pattern for DeMo when extended to two nodes, with inter-node communication. DeMo does not employ any sharding, and is thus using an all_gather operation across the nodes, in the case of multiple nodes. This high amount of inter-node communication, is clearly evident in the figure.
Algorithm 1 FlexDeMo extended from DeMo
learning rate $\eta$, decay $\beta\in(0,1)$, parameters $\theta_{t}$, momentum $m_{t}$,
sharding-set $S$, replication-set $R$, hyperparameters $s,k$
$\theta_{t}^{i}\leftarrow GradReduceScatter(\theta_{t},S)$ $\triangleright$ Get local parameter shard. Intra-Node.
$\Delta_{t}^{i}\leftarrow LocalSGD(\theta_{t}^{i})$ $\triangleright$ Get local gradient $\Delta_{t}$
$m_{t}\leftarrow\beta m_{t}+\Delta_{t}$ $\triangleright$ Accumulate gradient in momentum $m$
$q_{t}\leftarrow ExtractFastComponents(m_{t},s,k)$ $\triangleright$ Extract fast components $q$ from $m$
$m_{t+1}\leftarrow m_{t}-q_{t}$ $\triangleright$ Remove q from m
$Q_{t}^{i}\leftarrow Synchronize(q_{t})$ $Q_{t}^{i}\leftarrow Synchronize(q_{t},R)$ $\triangleright$ Synchronize across all nodes. Inter-Node.
$\theta_{t+1}^{i}\leftarrow\theta_{t}^{i}-\eta Q_{t}$ $\triangleright$ Parameter update step
FlexDeMo degrades gracefully to pure FSDP and pure DDP settings for trivial sharding and replication groups: In the edge-case of only sharding ($\left|R\right|$ = 1), the behaviour of FlexDeMo corresponds to FSDP. In the edge-case of only replication ($\left|S\right|$ = 1), the behavior of FlexDeMo corresponds to DDP with DeMo-style replication. Without sharding and replication, the behavior of FlexDeMo collapses to single-accelerator training with the underlying optimizer.
Replication Schemes
We are introducing the Random and Striding replication schemes and compare with the already proposed DiLoCo. Random denotes Random selection of indices. Striding denotes the selection of the n’th index and DiLoCo denotes synchronization each n’th optimization step. These replicators do not operate in chunks as the DeMo replicator. All replication schemes can achieve the same relative data compression. However, Striding and Random schemes do not need to share the selected indices with the gradient, as their operations can reproduced across training-instances with a fixed seed, enabling us to share double the amount of data, on the same bandwidth.
Decoupled AdamW
AdamW is defacto the default optimizer for most objectives, prompting us to believe is also would demonstrate strong performance with these replication schemes. We implement AdamW, with reducing communication in mind, hence we do not share the first and seconds momenta, which would require 2-3 times more communication. The exponential moving average (EMA) and the moving average of the squared gradients are therefore not synchronized, and neither are the max of all exponential moving averages of squared gradients, if amsmax is employed.
Experiments
We present a series of experiments on FlexDeMo covering encoder-decoder language models, vision transformers, and decoder-only language models including scaling experiments, demonstraing the capabilties and training-behaviour across domains. We compare our Decoupled AdamW to DeMo-SGD as the underlying optimizer, challenging the DeMo replication-scheme with our introduced Striding and Random schemes, and further, compare with DiLoCo. Last, we demonstrate the performance and efficiency comparing performance to standard Hybrid-FSDP with a standard AdamW optimizer (Kingma and Ba 2015; Loshchilov and Hutter 2019) demonstrating the effectiveness of FlexDeMo.
Replication Schemes
Setup
First, we study hyperparameters for DeMo replication, including sign, topKK, transfer-dtype-type. Experiments can be found in the Appendix B. Here, we focus on small-scale experiments to study method design choices and hyperparameters. Specifically, we investigate the characteristics of the decoupled optimizations scheme in encoder-decoder architecture by training with a sequence-to-sequence objective. We train the T5-Large model on its native translation task on the English-to-French subset of the Opus Books Dataset (Tiedemann 2012).
The experiments are conducted on 2 nodes with 2 accelerators, employing Hybrid-FSDP. We run for 20 epochs, with a learning rate of $1e^{-3}$ and no weight decay. The optimizer and the replication scheme is depending on the specific experiment. These experiments are carried out on a HPC with a dragonfly network architecture with 200 Gbps interconnect. Nodes are NUMA-nodes and comprises of two GPU modules each with 4 AMD MI250x GPUs (128GB per module). The two modules are connected with infinity fabric connections (50+50 Gb/s) and the modules communicate over the network. We base our foundation experiments on the Opus Books dataset subset of English to French, standardizing our experimental setup for reproducibility.
Results: SGD vs. Decoupled AdamW
AdamW (Kingma and Ba 2015) is de-facto default optimizer, having demonstrated superior performance across many domains and tasks. However DeMo (Peng, Quesnelle, and Kingma 2024) is employing SGD (which we denote DeMo-SGD, we differentiate as it accumulates momenta.) as underlying optimizer. To investigate this design-choice and broaden the investigation, we set out to compare DeMo-SGD and our Decoupled-AdamW with employing DeMo, DiLoCo (Douillard et al. 2023), Random, and Striding replication schemes. We report validation loss in Figure 1. The experiments are conducted such that the bandwidth usage is identical in each setting (we denote this as the compression-rate), as it would be defined by hardware in practice. DeMo-SGD with Random replication demonstrates best performance. DeMo-SGD yields clearly better at generalization to the validation-set, with Random replication being best. Decoupled-AdamW and DeMo-SGD with DiLoCo replication following as seconds and third best, respectively. DeMo-SGD is improving overall as fast as the DiLoCo variants, but overall worse. Decoupled-AdamW with DeMo and Random replication schemes yields worst performance.
Figure 1: T5-Large OpusBooks English-French, validation loss for DeMo SGD and Decoupled AdamW. The compression rate is chosen to keep the bandwidth usage constant across replication methods. DeMo-SGD with Random replication is clearly superior converging faster than both DiLoCo variants and DeMo SGD with DeMo replicator. Decoupled AdamW behaves similarly to SGD-DeMo with DiLoCo replicator, with an offset on the y-axis
T5: Translation Task
Setup
We train a T5-base model on the French-to-English subset of Opus Books dataset (Tiedemann 2012) for 20 epochs, in a Hybrid-FSDP setup with 2 nodes with 8 accelerators each. We employ the SGD optimizer with Random, DeMo, DiLoCo and Striding replication. DeMo employs a chunk-size of 32. We hold out 20% as validation data.
Results
Figure 2(a) shows the validation loss. Training loss is available in Appendix C. Random replication performs best, with compressions $\frac{1}{2}$ and $\frac{1}{4}$ being best, respectively. Subsequently, DeMo $\frac{1}{8}$ comes followed by DeMo $\frac{1}{4}$, $\frac{1}{16}$ and $\frac{1}{32}$. DiLoCo and Striding replication schemes generally convergences much slower. We see similar results for Random $\frac{1}{16}$ and $\frac{1}{32}$, where the signal seems to become too sparse, yet still learning.
Figure 2: Epoch train and validation loss for T5-Large on the Opus Books En-Fr subset.
(a) T5-Large Validation loss on the Opus Books En-Fr subset. Random and DeMo replication demonstrates strong performance.
(b) ViT-B Validation loss on Cifar100. DeMo and DiLoCo performs the best and very comparable.
Discussion
We have shown that DeMo-SGD with Random-replication schemes is best for T5 of the replication schemes, with compression rates $\frac{1}{2}$ and $\frac{1}{4}$, whereas DeMo lands at second place, with compression rates $\frac{1}{8}$ and $\frac{1}{4}$, respectively. While DeMo enables more compression in its best-performing configurations, Random yields lower loss and is faster in practice. Striding and DiLoCo are converging slowly, making them less competitive here.
ViT: Image Classification
Setup
We train a ViT-B (Patch $16$, 224224x224224) model for vision classification on the Cifar100 dataset (Krizhevsky, Hinton et al. 2009), for 20 epochs, in a Hybrid-FSDP setup with $2$ nodes with $8$ accelerators each. experiments are run with DeMo-SDG as the underlying optimizer with a learning rate of $1e^{-5}$.
Results
We report our findings in Figure 2(b). Training loss is available in the Appendix D. DeMo replication with a compression rate of $\frac{1}{2}$ and $\frac{1}{4}$ demonstrates the highest performance (comparable), closely followed by DeMo at $\frac{1}{16}$, with DiLoCo $\frac{1}{2}$ in-between the two. We generally see that the Random scheme struggles quite clearly in this domain. Striding replication even performs better than Random.
Discussion
Contrary to the encoder-decoder architecture on a next-token-prediction training with superior Random replication scheme, DeMo replication is superior here. We hypothesis that this is because fast moving momenta is more suited for this task, as less pronounced features doesn’t contribute to the learning-objective, potentially inducing noise, whereas these small nuances in language modeling is very valuable. Randomly chosen indices can include feature-information not important for the implicit supervised contrastive learning. DiLoCo with a quite low compression follows DeMo’s performance closely, while taking a hit on the performance when increasing the compression rate. Striding does follow DiLoCo’s performance closely, we hypothesis that the structure provided by this scheme work in highly structured data, that images are.
OLMo2: Causal Language Modeling
Setup
We train the OLMo2 decoder-only model for causal language modeling, in a Hybrid-FSDP setup with $2$ nodes with $4$ accelerators each for 1B parameters. We train on the Dolma v1.6 dataset (Soldaini et al. 2024) for 10K steps. We use the OLMo’s standard parameters [^1] https://github.com/allenai/OLMo/blob/main/configs/official-0425/OLMo2-1B-stage1.yaml, with the chance of 4% warm-up steps. Each node is composed by a custom Ampere A100 64GB GPU, with 2x dual-port HDR network interface (400Gbps aggregated). The nodes are connected in a infiniband-based Dragonfly+ topology with a low latency interconnect with 200Gb/s. The purpose of our experiments is to demonstrate FlexDeMo’s reduction in communication overhead in comparison with and Hybrid-FSDP and a conventional AdamW optimizer (Kingma and Ba 2015; Loshchilov and Hutter 2019).
Results
We report our findings in Figures 3 and 4. All of the FlexDeMo with DeMo replication demonstrates the best performance, with interestingly the compression-rate $\frac{1}{32}$, being the best, closely followed by $\frac{1}{16}$, demonstrating some interesting artifacts of high compression. Random-replication is next-best, and better than DeMo with compression $\frac{1}{4}$, which performs comparable to Random $\frac{1}{16}$. Hybrid-FSDP with conventional AdamW (red), performs better all DiLoCo configurations, but worse than all DeMo and Random configurations. Strided replication generally yields unstable training and uncompetitive results. In Figure 4, we see that all replicator-types across all compression-rates is significantly faster than conventional AdamW, which employs full replication. Conducting timings are quite unreliable on a HPC, as it depends on the network-congestion during the time of the experiment, preventing us to make exact comparisons between the replicators.
Figure 3: OLMo2 1B train loss (zoomed) over 10K training steps on Dolma v1.6 using different replicators and compression rates. All experiments, except the Hybrid-FSDP baseline with AdamW on 2 nodes, use DeMo-SGD.
Figure 4: OLMo2 1B train loss vs. wall-clock time over 10K training steps on Dolma v1.6, comparing different replicators and compression rates. All experiments use DeMo-SGD except for the Hybrid-FSDP baseline with AdamW. FlexDeMo shows clear improvements in convergence speed.
Discussion
Similarity, as with the ViT-archtiecture the DeMo-replicator performs best for OLMo-architecture with compression rates $\frac{1}{32}$ and $\frac{1}{16}$, being the best. This is somewhat closely followed the Random replication scheme, however with larger compressions on $\frac{1}{4},\frac{1}{8}$ and $\frac{1}{2}$, respectively, making them even less attractive to the DeMo variant. Interstingly, we do not see an optimal replication-scheme within each We demonstrated the impact of FlexDeMo with all the replicators compared to conventional Hybrid-FSDP and AdamW across only 2 nodes on a HPC. We see that all the replicator schemes yields training-times around approx. 2.6 times faster than the Hybrid-FSDP with AdamW setting. As timing on HPC systems is dependent on the current network congestion, we refrain from commenting on differences between the replicators, as they probably will alternate when reproducing. However, this does not make the big relative margin to the conventional setup uncertain and paves the way for better utilizing GPU allocations and lowering the environmental impact significantly. Thus, with the hybrid-sharding strategy, the all_gather is done only once per node, instead of once per accelerator in the original decoupled momentum implementation, which is clearly beneficial. With introducing hybrid-sharding, we need to use the gather-operation intra-node, which is reasonably fast, but can be optimized by implementing hooks and streams, which we leave for future work.
OLMo2: Scaling Up
Setup
In this experiment we investigate the performance and efficiency of FlexDeMo at scale. Our setup is based the experiments in Figure 3, employing the Demo-SGD optimizer, and both DeMo and Random-replication schemes with a $\frac{1}{32}$ compression rate and a learning-rate of $1e^{-3}$.
Results
We report our findings in Figure 5. Similarly, we see that Random replication performs worse, than DeMo replication, which performs almost comparable to conventional AdamW (full-sync). However, DeMo is significantly slower. We attribute the to DeMo’s dependence on the all_gather operation, which prevents is to scale meaningfully, shown In Figure 6. The values on the $y$-axis is smaller than in 3, due to only tracking on rank 0, to not induce a bottleneck for tracking.
Discussion
In Figures 5 and 6, we investigate the scale both Random and DeMo replication schemes, powered by the DeMo-SGD optimizer and comparing against conventional Hybrid-FSDP with AdamW optimization. While DeMo replications performance is clearly best, there is a big drawback in the time-scaling of the scheme, due to the all_gather operation not scaling, as one would expect, due to the blocking nature of the operation. However, Random is around 64% faster than the conventional setup. This could call for 2-stage distributed training setup, employing Random replication for the majority of the training, employing the conventional setup for a subsequent stage.
Figure 5: OLMo2 1B train loss (zoomed) over 10K training steps on Dolma v1.6 using DeMo-SGD with a 1/32 compression rate on 64 nodes. We compare repl. schemes DeMo and Random against the Hybrid-FSDP + AdamW baseline.
Figure 6: OLMo2 1B train loss vs. wall-clock time (in minutes) on 64 nodes over 10K training steps using DeMo-SGD. DeMo does not scale efficiently with node count due to overhead from the all_gather operation, especially compared to Hybrid-FSDP with AdamW.
General Discussion and Conclusion
Through a set of experiments, we have validated important hyper-parameters, investigating the characteristics of FlexDemo, with DeMo replication, including communication data type, underlying optimizer, the question on the sign function and TopKK. We do expect parameters to be architecture and potentially domain-dependent.
We demonstrate that using sign before synchronizing is a corner-stone in this optimization scheme. Immediately, this tells us that direction is more important than magnitude for our case of decoupled optimization. This has also been studied in prior work (Bernstein et al. 2018). The ternary system opens up for the possibility to compress the data even more, assuming it can be done fast enough.
In experiments with full control over the network bandwidth between the nodes, we have shown how Random replication scales across different replicators and compression rates . This, with the DeMo-SGD optimizer, is not surprising, as SGD employs fewer parameters per model parameter than AdamW.
We have introduced a decoupled variant of AdamW, investigating a version not sharing the first and second moments. However, we demonstrate in Figure 1, that Decoupled-AdamW is not generally superior and only outperforms DeMo-SGD when compared on the DiLoCo-scheme. We did not develop this optimizer further, as we in OLMo2 demonstrated superior performance relative to conventional Hybrid-FSDP with AdamW.
We have studied the Striding and Random replication schemes we introduced to facilitate FlexDeMo and challenged experimental choices made in DeMo. Using any of these two schemes enables us to utilize more of the bandwidth sending actual data and not the corresponding indices. We demonstrate that DeMo works best on the ViT and decoder-architectures, whereas Random performs best in the encoder-decoders architecture calling for investigation into more architecture-specific replications.
Throughout our experiments we have experimented with both NCCL and RCCL communications-backends, demonstrating DeToNATION and FlexDeMo across both platforms. Overall, this opens an avenue of future research of decoupled training-techniques, extending these schemes but also, importantly, investigating sparse optimization in different domains. This is important, as we demonstrated that different replicators are optimal for different tasks. For future work, we leave an open question of weather the type of replicator depends on the task or architecture.
Future Work
Introducing FSDP requires the gathering of full parameters, hence an initial reduce_scatter-operation on each gradient. PyTorch’s FSDP implementation does this when the gradients become available during the backward-pass, which we have disables with no_sync. However, this can be implemented using hooks, which will enable starting already during the backward-pass. Further, we believe that improving communication with CUDA streams in a general asynchronous synchronization scheme contains a substantial efficiency potential. Introducing FSDP2222https://docs.pytorch.org/docs/stable/
distributed.fsdp.fully_shard.html#pytorch-fsdp2-fully-shard
or SimpleFSDP (Zhang et al. 2024) will also speed up the basic communication operations. These implementations will make the groundwork relaxing the memory-requirements even further, more comfortably enabling cross-node sharding and, thus, supporting very large models. Further, this opens for the feasibility of connecting one or more HPC systems when training, simplifying collaborations and utilizing the existing resources.
Conclusion
In conclusion, we have shown that FlexDeMo successfully extends DeMo optimization to the setting of hybrid sharded FSDP. Our experiments indicate results better than AdamW, while at the same time being significantly faster, demonstrating the potential to reduce the bandwidth needed for communication between nodes. We can conclude that DeMo-SGD proves to be the superior underlying optimizer in most setups, while the underlying replicator is more dependent on given architecture and task. We also introduced Decoupled AdamW. DeMo replication scales significantly worse than the conventional scheme, whereas Random, performing a little worse, keeps delivering across many nodes. Overall, this enables FlexDeMo to be used to optimize LLMs that do not fit within the memory of one accelerator and further supports training LLMs in low-resource network settings, ultimately lowering the entry point for both practitioners and researchers to participate in the development of future models and use cases, while offering more compute from the same GPU hours, potentially reducing the overall environmental impact.
Appendix A A: Communication DeMo vs. FlexDeMo
In Figure 7 we give a detailed visualization of the communication in DeMo and FlexDeMo using hybrid-sharding.
(a) DeMo - 1 Node
(b) DeMo - 2 Nodes
(c) FlexDeMo - 2 Nodes
Figure 7: Communication Patterns of DeMo and FlexDeMo, with their respective intra-node and inter-node communication.
Appendix B B: DeMo Hyperparameters
TopK
TopK hyperparameter, defines the number of fastest moving momenta which is synchronized between replications, when using the DeMo replicator. We conduct an experiment, investigating the optimal TopK, for our setup. We report our results in Figure 8. Top44 is clearly superior for our setting, followed by Top22, Top11, Top88, respectively. Last, with a huge degradation in performance Top1616.
Figure 8: T5-Large on OpusBooks En-Fr over different TopK’s. Demo-SGD optimizer and DeMo-replicator.
Sign vs. No Sign.
The sign-function [^3] https://docs.pytorch.org/docs/2.7/generated/torch.sign.html introduces great reduction in required data transferred, reducing the actual communication to a ternary system with minimal information exchange. This design-choice was introduced in DeMo (Peng, Quesnelle, and Kingma 2024), however, not justified. We report loss on the validation set in Figure 9. Interestingly, employing the sign-function to gradients before synchronization, generally yields significant performance-gains with a huge gap in the validation-results with DeMo and DiLoCo leading performance, closely followed by the Random scheme, with Striding under-performing quite significantly.
Figure 9: Validation loss evaluations on T5-Large trained on OpusBooks (En-Fr) demonstrate the impact of employing the sign function prior to gradient synchronization. Very clearly sign have a general positive impact, especially for the sparse optimization schemes.
Average Time per Optimization Step
We analyze the average time per step for $10$, $100$, $1000$ and 10.00010{.}000 Mbps network connections training the T5-Large and ViT-B. The experiment is conducted on a controlled network between two local dedicated nodes. We compare DeMo and Random replication schemes with underlying DeMo-SGD optimizer with compression-rates $\frac{1}{16}$ and $\frac{1}{32}$, against the Decoupled-AdamW optimizer (ours). We report our results in Figure 10. As expected, as the bandwidth decreases, the importance of the compression-rate becomes more pronounced, where the compression of $\frac{1}{32}$ clearly demonstrates to be fastest followed by $\frac{1}{16}$ with $\frac{1}{1}$, being last, dictated by the amount of data to communicate, with Random DeMo-SGD being the fastest, as demonstrated both in Figure 10(a) and 10(b). This is especially significant at bandwidths lower than 500Mbps, which is a level of bandwidths comparable to those available in HPCs, in practice. Random replication is approx. $3.33$ times faster than DeMo-replication at 1010Mbps, while being approx. $18$ times faster than Decoupled-AdamW with full replication (Figure 10(b)). Generally we see that Random replication with compression $\frac{1}{16}$ scales similarly to DeMo replication with compression $\frac{1}{32}$, demonstrating exactly that Random replication is around twice as fast as DeMo, as expected from DeMo transferring twice the amount of data, at the same compression rate.
(a) T5-Large on 2 nodes across employing different replicators using DeMo-SGD and Decoupled AdamW as base optimizers
(b) ViT-B on 2 nodes across employing different replicators using DeMo-SGD and Decoupled AdamW as base optimizers
Figure 10: Average time per optimizer step for T5-Large and ViT-B in bandwidth-constrained setups.
DeMo Chunk Size
We conduct experiments with the DeMo-replicator across different chunk-sizes with compression rates of $1/16$ and $1/8$, investigating the training behavior and performance yields. We report our findings in Figure 11, with the corresponding bandwidth usage in Figure 12. A compression rate of $1/8$ and the smallest chunk-size of 16 demonstrates slightly better performance, generally getting worse as the chunk-size increases. Interestingly, we see the opposite tendency for a compression rate $1/16$, where the larger chunk sizes yields better performances. Compression rate $1/16$ with chunk-size 96 outperforms chunk-size $128$ with compression-rate $1/8$. However, this is not a general tendency.
Figure 11: SGD DeMo Validation Loss Curve for Compression Rates $1/16$ and $1/8$ over the chunk-sizes $16,32,64,96,18,192$ and $256$.
Figure 12: Bandwidth usage for chunk size variations.
Communication Floating Point Data Type (dtype)
The most significant momenta are shared in full precision by previous work (Peng, Quesnelle, and Kingma 2024). Naturally, we need to confirm the necessity to transfer double precision on the cost of doubled data footprint, visualized in Figure 13. We validate this choice in Figure 14, where we confirm DeMo and full precision is superior, followed close by full sync (DiLoCo) full-precision. Interestingly, the full-sync scheme is mostly unaffected by the data precision, whereas both DeMo and Random replication change significantly with precision, where higher precision, leads to better performance.
Figure 13: Bandwidth usage for each precision-type.
Figure 14: Validation loss for different precision data types.
Appendix C C: T5-Large Train Losses
In Figure 15 we report the train-losses for the T5 translation task experiments in the main paper.
Figure 15: T5-Large Train loss on the Opus Books En-Fr subset.
Appendix D D: ViT Train Losses
In Figure 16 we report the train-losses for the ViT image classification experiments in the main paper.
Figure 16: Epoch train loss for Vit-B 224x224 on Cifar100
Appendix E E: Ethical Considerations
Our work improves efficiency of distributed training. It may contribute to the democratization of training large language models and to reducing the environmental footprint of LLMs. In particular, our findings suggest that LLMs potentially can be trained on a lower budget. We acknowledge that this may make training such models more accessible. However, this work is purely scientific and does not promote easier access to systems for training such models. Using our proposed method, it might be possible to reduce compute budgets, making larger models more financially accessible.
Appendix F Reproducibility
[…]
Note that some HPCs are restricted in terms of downloads from the web on the compute nodes, and the models and datasets should thus be download accordingly to each setup. Scripts are included to download the datasets, where relevant.